OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

Use point-slope form: y-y1=m(x-x1), where m=2 and y1=5, x1=-1. y-5 = 2(x+1) Write answer in this form, Standard Form: ax+by=c Can you do this and tell me what you got?